OpenStudy (anonymous):

Please help I have no idea what to do! Find an equation for the nth term of the arithmetic sequence. a19 = -58, a21 = -164

OpenStudy (anonymous):

So the formula for the nth term in an arithmetic sequence is: Un = a + d(n-1)

OpenStudy (anonymous):

Where n = the value of the term you are trying to find.

OpenStudy (anonymous):

so I solve for n?

OpenStudy (anonymous):

No, you find the difference between each term (d in the formula above) and the first term of the arithmetic sequence (a in the formula above)

OpenStudy (anonymous):

so -58+164? finding the difference?

OpenStudy (anonymous):

The common difference is (-164+58)/2. Which equals -53

OpenStudy (anonymous):

So the answer is an = 896 - 53(n - 1) ?

OpenStudy (anonymous):

Yes and your teacher/tutor might expect you to expand those brackets to make it: Un = 896 -53n +53 = 949 - 53n

OpenStudy (anonymous):

But both would be correct.

OpenStudy (anonymous):

you are great!!!!

OpenStudy (anonymous):

:)